# A convergence proof for a finite element discretization of Chorin’s projection method of the incompressible Navier-Stokes equations

### v1

@@ 3.4. Energy stability
$\frac{1}{2\Delta t}\left(\left\lVert\widetilde{u}^{m+1}_{h}\right\rVert_{L^{2}}^{2}-\left\lVert u^{m}_{h}\right\rVert^{2}_{L^{2}}+\left\lVert\widetilde{u}_{h}^{m+1}-u^{m}_{h}\right\rVert_{L^{2}}^{2}\right)+b(u^{m}_{h},\widetilde{u}^{m+1}_{h},\widetilde{u}_{h}^{m+1})\\ =-\mu(\nabla\widetilde{u}_{h}^{m+1},\nabla\widetilde{u}_{h}^{m+1})+(\mathcal{P}_{\mathbb{U}_{h}}f^{m},\widetilde{u}^{m+1}_{h}).$
@@
This looks pretty good, except we would prefer to have $u_{h}^{m+1}$ instead of $\widetilde{u}^{m+1}_{h}$ in the kinetic part of the energy. Therefore, we take $v=u^{m+1}_{h}\in\mathbb{U}_{h}$ as a test function in (3.12a). Then since it satisfies (3.12b) for all $q\in\mathbb{P}_{h}$ and $p^{m+1}_{h}\in\mathbb{P}_{h}$, we see that the second term is zero and hence
@@
where we have used the identity $2a(a-b)=(a^{2}-b^{2}+(a-b)^{2})$ once more. Using this identity in (3.19), and the property of the $L^{2}$-projection, (3.6), we obtain
@@ 4. Convergence analysis of the scheme
(4.2) $\bar{u}_{h}(t)$ $=u_{h}^{m},\qquad$ $t\in[t^{m},t^{m+1}),$
@@
From the energy balance in the last section, Lemma 3.5 and Remark 3.6, we obtain the following uniform a priori estimates for the sequences $\{u_{h}\}_{h>0}$, $\{\bar{u}_{h}\}_{h>0}$, and $\{\widetilde{u}_{h}\}_{h>0}$:
$\{u_{h}\}_{h>0},\{\bar{u}_{h}\}_{h>0}\subset L^{\infty}(0,T;L^{2}_{\text{div}}(\Omega)),$
@@
(4.6) $u_{h}\overset{\star}{\rightharpoonup}u,$ $\quad\text{in }\,L^{\infty}(0,T;L^{2}_{\text{div}}(\Omega)),$
(4.7) $\bar{u}_{h}\overset{\star}{\rightharpoonup}\bar{u},$ $\quad\text{in }\,L^{\infty}(0,T;L^{2}_{\text{div}}(\Omega)),$
(4.8) $\widetilde{u}_{h}\overset{\star}{\rightharpoonup}\widetilde{u}$, $\quad\text{in }\,L^{\infty}(0,T;L^{2}(\Omega)),$
In order to prove convergence of the scheme, we will need to derive strong precompactness of the sequences in $L^{2}([0,T]\times\Omega)$. To achieve this, we will use a version of the Aubin-Lions-Simon lemma [30, 27]. This requires a uniform bound on the time derivative of $u_{h}$. We prove this in the following lemma:
@@
Under the assumption that $h\leq C\sqrt{\Delta t}$, the time derivative of $u_{h}$ satisfies uniformly in $h>0$:
(4.10) $\{\partial_{t}u_{h}\}_{h>0}\subset L^{2}(0,T;(H_{\operatorname{div}}^{1}(\Omega)\cap H^{2}(\Omega))^{*}).$
@@
$\partial_{t}u_{h}=\frac{u^{m+1}_{h}-u^{m}_{h}}{\Delta t},\quad t\in[t^{m},t^{m+1}).$
Recall that $\mathcal{P}_{\mathbb{U}_{h}}:L^{2}(\Omega)^{d}\to\mathbb{U}_{h}$ is the $L^{2}$-orthogonal projection onto the finite element space. Then we have for $v\in L^{2}(0,T;H^{1}_{\operatorname{div}}\cap H^{2}(\Omega))$, using (3.11) and (3.12a)
$\int_{0}^{T}(\partial_{t}u_{h},v)dt$ $=\int_{0}^{T}(\partial_{t}u_{h},v-\mathcal{P}_{\mathbb{U}_{h}}v)dt+\int_{0}^{T}(\partial_{t}u_{h},\mathcal{P}_{\mathbb{U}_{h}}v)dt$
$=\underbrace{\int_{0}^{T}(\partial_{t}u_{h},v-\mathcal{P}_{\mathbb{U}_{h}}v)dt}_{\text{I}}-\underbrace{\int_{0}^{T}b(\bar{u}_{h},\widetilde{u}_{h},\mathcal{P}_{\mathbb{U}_{h}}v)dt}_{\text{II}}-\underbrace{\int_{0}^{T}\mu(\nabla\widetilde{u}_{h},\nabla\mathcal{P}_{\mathbb{U}_{h}}v)dt}_{\text{III}}$
$\quad+\underbrace{\int_{0}^{T}(p_{h},\operatorname{div}(\mathcal{P}_{\mathbb{U}_{h}}v-v))dt}_{\text{IV}}+\underbrace{\int_{0}^{T}(f_{h},\mathcal{P}_{\mathbb{U}_{h}}v)dt}_{\text{V}},$
where we have used that $v$ is divergence free. We estimate each of the five terms. Since $\partial_{t}u_{h}\in\mathbb{U}_{h}$, we have that
$(\partial_{t}u_{h},v)=(\partial_{t}u_{h},\mathcal{P}_{\mathbb{U}_{h}}v),$
@@
$|\text{II}|$ $\leq\left|\sum_{m=0}^{N-1}\int_{t^{m}}^{t^{m+1}}b(u_{h}^{m},\widetilde{u}_{h}^{m+1},\mathcal{P}_{\mathbb{U}_{h}}v)dt\right|$
$\leq\sum_{m=0}^{N-1}\left\lVert u^{m}_{h}\right\rVert_{L^{2}}\left\lVert\nabla\widetilde{u}_{h}^{m+1}\right\rVert_{L^{2}}\int_{t^{m}}^{t^{m+1}}\left\lVert\mathcal{P}_{\mathbb{U}_{h}}v\right\rVert_{L^{\infty}}dt$
$\leq\max_{k=0,\dots,N-1}\left\lVert u^{k}_{h}\right\rVert_{L^{2}}\left(\Delta t\sum_{m=0}^{N-1}\left\lVert\nabla\widetilde{u}_{h}^{m+1}\right\rVert_{L^{2}}^{2}\right)^{\frac{1}{2}}\left\lVert\mathcal{P}_{\mathbb{U}_{h}}v\right\rVert_{L^{2}(0,T;L^{\infty}(\Omega))}$
@@
where we used Morrey’s inequality, the Sobolev embedding theorem and the stability of the $L^{2}$-projection operator, equation (3.7a) (see e.g. [8]), for the last inequalities. We proceed to term III:
$|\text{III}|$ $\leq\mu\left|\sum_{m=0}^{N-1}\int_{t^{m}}^{t^{m+1}}(\nabla\widetilde{u}^{m+1}_{h},\nabla\mathcal{P}_{\mathbb{U}_{h}}v)dt\right|$
$\leq\mu\left(\Delta t\sum_{m=0}^{N-1}\left\lVert\nabla\widetilde{u}^{m+1}_{h}\right\rVert_{L^{2}}^{2}\right)^{\frac{1}{2}}\left\lVert\nabla\mathcal{P}_{\mathbb{U}_{h}}v\right\rVert_{L^{2}([0,T]\times\Omega)}$
@@
$\leq\frac{C(E_{h}^{0}+\left\lVert f\right\rVert_{L^{2}([0,T]\times\Omega)}^{2})^{1/2}h}{\sqrt{\Delta t}}\left\lVert v\right\rVert_{L^{2}(0,T;H^{2}(\Omega))},$
@@
$|\text{V}|\leq\left|\int_{0}^{T}(f_{h},v)dt\right|\leq\left\lVert f_{h}\right\rVert_{L^{2}(0,T;L^{2}(\Omega))}\left\lVert v\right\rVert_{L^{2}(0,T;L^{2}(\Omega))}\leq C\left\lVert f\right\rVert_{L^{2}(0,T;L^{2}(\Omega))}\left\lVert v\right\rVert_{L^{2}(0,T;L^{2}(\Omega))}.$
@@
$\sup_{0\neq v\in L^{2}(0,T;H^{1}_{\operatorname{div}}\cap H^{2})}\left|\frac{(\partial_{t}u_{h},v)}{\left\lVert v\right\rVert_{L^{2}(0,T;H^{1}_{\operatorname{div}}\cap H^{2})}}\right|\leq C\left(1+\frac{h}{\sqrt{\Delta t}}\right),$
which is uniformly bounded under the assumption that $h\leq C\sqrt{\Delta t}$. Thus $\partial_{t}u_{h}\subset L^{2}(0,T;(H^{1}_{\operatorname{div}}(\Omega)\cap H^{2}(\Omega))^{*})$ uniformly in $h>0$. ∎
Next, we will show that the sequences $\{\widetilde{u}_{h}\}_{h>0}$, $\{u_{h}\}_{h>0}$, $\{\bar{u}_{h}\}_{h>0}$ have the same weak limits $\widetilde{u}=u=\bar{u}$:
@@
Assume that $\Delta t=o_{h\to 0}(1)$. Then, we have $u=\bar{u}=\widetilde{u}$ a.e. in $0,T]\times\Omega$ and in $L^{2}([0,T]\times\Omega)$, and
(4.12) $\lim_{h\to 0}\left\lVert u_{h}-\bar{u}_{h}\right\rVert_{L^{2}([0,T]\times\Omega)}$ $=0,$
@@
$\left\lVert\bar{u}_{h}-\widetilde{u}_{h}\right\rVert_{L^{2}([0,T]\times\Omega)}^{2}$ $=\Delta t\sum_{m=0}^{N-1}\int_{\Omega}\left|u^{m}_{h}-\widetilde{u}^{m+1}_{h}\right|^{2}dx$
@@
by the energy balance, Lemma 3.5 and Remark 3.6. This shows (4.13). Using this, it then follows easily that for any test function $\varphi\in L^{2}([0,T]\times\Omega)$,
@@
$\left\lVert\bar{u}_{h}-u_{h}\right\rVert_{L^{2}([0,T]\times\Omega)}^{2}$ $=\sum_{m=0}^{N-1}\int_{t^{m}}^{t^{m+1}}\!\!\int_{\Omega}\left|\frac{t-t^{m}}{\Delta t}(u^{m+1}_{h}-u^{m}_{h})\right|^{2}dxdt$
$\leq\Delta t\sum_{m=0}^{N-1}\int_{\Omega}\left|u^{m+1}_{h}-u^{m}_{h}\right|^{2}dx$
$\leq 2\Delta t\sum_{m=0}^{N-1}\int_{\Omega}\left|u^{m+1}_{h}-\widetilde{u}^{m+1}_{h}\right|^{2}+\left|\widetilde{u}^{m+1}_{h}-u^{m}_{h}\right|^{2}dx$
@@
Thus as $\Delta t,h\to 0$, we obtain (4.12), and in the same way as before, one shows that the limits $u$ and $\bar{u}$ agree almost everywhere. ∎
Next, we would like to apply the Aubin-Lions-Simon lemma [27] and obtain a strongly convergent subsequence of $\{u_{h}\}_{h>0}$ in $L^{2}([0,T]\times\Omega)$. However, we have on one hand the time continuity for $u_{h}$ (with values in the dual space of $H^{1}_{\operatorname{div}}\cap H^{2}$) and on the other hand spatial regularity in $H^{1}_{0}(\Omega)$ for the sequence of approximations $\{\widetilde{u}_{h}\}_{h>0}$. Since $\{\widetilde{u}_{h}\}_{h>0}$ are not necessarily divergence free, in order to apply the Aubin-Lions-Simon lemma, we would need time continuity with values in a space that $L^{2}(\Omega)$ embeds into continuously. However, this may not be possible to achieve due to the lack of estimates for the pressure variable $p_{h}$. On the other hand, we do not have sufficient spatial regularity, e.g. $H^{1}$-bounds, for the variable $u_{h}$ from the energy inequality. When the domain is smooth enough, $C^{1,1}$, or convex, then the Leray projection is a continuous operator from $H^{1}$ to $H^{1}$ (see e.g. [4, Proposition IV.3.7.][7][24]). However, in a typical practical setting, the domain may be polygonal and nonconvex and so this assumption may not be satisfied. We will therefore need a variant of the Aubin-Lions-Simon lemma adapted to our setting. To prove it, the following classical lemma by Lions [30] will be useful:
@@
[30, Ch.III, Lem. 2.1] Let $X\subset Y\subset Z$ be Banach spaces, such that the embedding $X\subset Y$ is compact and the embedding $Y\subset Z$ is continuous. Then for every $\eta>0$ there exists a constant $C_{\eta}$ depending on $\eta$ and the spaces $X,Y,Z$ such that
(4.14) $\left\lVert v\right\rVert_{Y}\leq\eta\left\lVert v\right\rVert_{X}+C_{\eta}\left\lVert v\right\rVert_{Z},\quad\forall\,v\in X.$

### v2

@@ 3.4. Energy stability
$\frac{1}{2\Delta t}\left(\left\lVert\widetilde{u}^{m+1}_{h}\right\rVert_{L^{2}}^{2}-\left\lVert u^{m}_{h}\right\rVert^{2}_{L^{2}}+\left\lVert\widetilde{u}_{h}^{m+1}-u^{m}_{h}\right\rVert_{L^{2}}^{2}\right)+b(\widetilde{u}^{m}_{h},\widetilde{u}^{m+1}_{h},\widetilde{u}_{h}^{m+1})\\ =-\mu(\nabla\widetilde{u}_{h}^{m+1},\nabla\widetilde{u}_{h}^{m+1})+(\mathcal{P}_{\mathbb{U}_{h}}f^{m},\widetilde{u}^{m+1}_{h}).$
@@
This looks promising, except we would prefer to have $u_{h}^{m+1}$ instead of $\widetilde{u}^{m+1}_{h}$ in the kinetic part of the energy. Therefore, we take $v=u^{m+1}_{h}\in\mathbb{U}_{h}$ as a test function in (3.12a). Then since $u^{m+1}_{h}$ satisfies (3.12b) for all $q\in\mathbb{P}_{h}$ and $p^{m+1}_{h}\in\mathbb{P}_{h}$, we see that the second term is zero and hence
@@
where we have used the identity $2a(a-b)=(a^{2}-b^{2}+(a-b)^{2})$ once more. Using this identity in (3.18), and the definition of the $L^{2}$-projection, (3.6), we obtain
@@ 4. Convergence analysis of the scheme
(4.2) $\bar{u}_{h}(t)$ $=\widetilde{u}_{h}^{m},\qquad$ $t\in[t^{m},t^{m+1}),$
(4.3) $\widehat{u}_{h}(t)$ $=\widetilde{u}_{h}^{m+1}+\frac{t-t^{m}}{\Delta t}(\widetilde{u}^{m+2}_{h}-\widetilde{u}^{m+1}_{h}),\qquad$ $t\in[t^{m},t^{m+1}),$
@@
From the energy balance in the last section, Lemma 3.4 and Remark 3.5, we obtain the following uniform a priori estimates for the sequences $\{u_{h}\}_{h>0}$, $\{\bar{u}_{h}\}_{h>0}$, $\{\widehat{u}_{h}\}_{h>0}$, and $\{\widetilde{u}_{h}\}_{h>0}$:
$\{u_{h}\}_{h>0}\subset L^{\infty}(0,T;L^{2}(\Omega)),$
$\{\bar{u}_{h}\}_{h>0},\{\widetilde{u}_{h}\}_{h>0},\{\widehat{u}_{h}\}_{h>0}\subset L^{\infty}(0,T;L^{2}(\Omega))\cap L^{2}(0,T;H^{1}_{0}(\Omega)).$
@@
(4.7) $u_{h}\overset{\star}{\rightharpoonup}u,\quad\bar{u}_{h}\overset{\star}{\rightharpoonup}\bar{u},\quad\widetilde{u}_{h}\overset{\star}{\rightharpoonup}\widetilde{u},\quad\widehat{u}_{h}\overset{\star}{\rightharpoonup}\widehat{u}$, $\quad\text{in }\,L^{\infty}(0,T;L^{2}(\Omega)),$
(4.8) $\bar{u}_{h}\rightharpoonup\bar{u},\quad\widetilde{u}_{h}\rightharpoonup\widetilde{u},\quad\widehat{u}_{h}\rightharpoonup\widehat{u}$, $\quad\text{in }\,L^{2}(0,T;H^{1}_{0}(\Omega)).$
In order to prove convergence of the scheme, we will need to derive strong precompactness of the sequences in $L^{2}([0,T]\times\Omega)$. To achieve this, we will need to adapt the Aubin-Lions-Simon lemma [29, 26] to our situation. This requires a uniform bound on the time derivative of $u_{h}$ or $\widehat{u}_{h}$. We prove this in the following lemma:
@@
Under the assumption that $h^{s-1}\leq C\sqrt{\Delta t}$, the time derivative of $\widehat{u}_{h}$ satisfies uniformly in $h>0$:
(4.9) $\{\partial_{t}\widehat{u}_{h}\}_{h>0}\subset L^{2}(0,T;(H_{\operatorname{div}}^{1}(\Omega)\cap H^{s}(\Omega))^{*}).$
for any $s\in\mathbb{N}$, $2\leq s\leq k+1$, where $k$ is the degree of the polynomial space used in $\mathbb{U}_{h}$.
@@
$\partial_{t}\widehat{u}_{h}=\frac{\widetilde{u}^{m+2}_{h}-\widetilde{u}_{h}^{m+1}}{\Delta t},\quad t\in(t^{m},t^{m+1}).$
Recall that $\mathcal{P}_{\mathbb{U}_{h}}:L^{2}(\Omega)^{d}\to\mathbb{U}_{h}$ is the $L^{2}$-orthogonal projection onto the finite element space. Then we have for $v\in L^{2}(0,T;H^{1}_{\operatorname{div}}\cap H^{s}(\Omega))$, using (3.11) and (3.12a)
$\int_{0}^{T}(\partial_{t}\widehat{u}_{h},v)dt$ $=\int_{0}^{T}(\partial_{t}\widehat{u}_{h},v-\mathcal{P}_{\mathbb{U}_{h}}v)dt+\int_{0}^{T}(\partial_{t}\widehat{u}_{h},\mathcal{P}_{\mathbb{U}_{h}}v)dt$
$=\underbrace{\int_{0}^{T}(\partial_{t}\widehat{u}_{h},v-\mathcal{P}_{\mathbb{U}_{h}}v)dt}_{\text{I}}-\underbrace{\int_{\Delta t}^{T+\Delta t}b(\bar{u}_{h},\widetilde{u}_{h},\mathcal{P}_{\mathbb{U}_{h}}v(t-\Delta t))dt}_{\text{II}}-\underbrace{\int_{\Delta t}^{T+\Delta t}\mu(\nabla\widetilde{u}_{h},\nabla\mathcal{P}_{\mathbb{U}_{h}}v(t-\Delta t))dt}_{\text{III}}$
$\quad+\underbrace{\int_{0}^{T}(p_{h},\operatorname{div}(\mathcal{P}_{\mathbb{U}_{h}}v-v))dt}_{\text{IV}}+\underbrace{\int_{\Delta t}^{T+\Delta t}(f_{h},\mathcal{P}_{\mathbb{U}_{h}}v(t-\Delta t))dt}_{\text{V}},$
where we have used that $v$ is divergence free. We extend $v$ and $f$ by zero for $t>T$. We estimate each of the five terms. Since $\partial_{t}\widehat{u}_{h}\in\mathbb{U}_{h}$, we have that
$(\partial_{t}\widehat{u}_{h},v)=(\partial_{t}\widehat{u}_{h},\mathcal{P}_{\mathbb{U}_{h}}v),$
@@
$|\text{II}|$ $\leq\left|\sum_{m=1}^{N}\int_{t^{m}}^{t^{m+1}}b(\widetilde{u}_{h}^{m},\widetilde{u}_{h}^{m+1},\mathcal{P}_{\mathbb{U}_{h}}v(t-\Delta t))dt\right|$
$\leq\sum_{m=1}^{N-1}\left(\left\lVert\widetilde{u}^{m}_{h}\right\rVert_{L^{2}}\left\lVert\nabla\widetilde{u}_{h}^{m+1}\right\rVert_{L^{2}}+\frac{1}{2}\left\lVert\widetilde{u}^{m+1}_{h}\right\rVert_{L^{2}}\left\lVert\operatorname{div}\widetilde{u}_{h}^{m}\right\rVert_{L^{2}}\right)\int_{t^{m}}^{t^{m+1}}\left\lVert\mathcal{P}_{\mathbb{U}_{h}}v(t-\Delta t)\right\rVert_{L^{\infty}}dt$
$\leq C\max_{k=1,\dots,N}\left\lVert\widetilde{u}^{k}_{h}\right\rVert_{L^{2}}\left(\Delta t\sum_{m=1}^{N}\left\lVert\nabla\widetilde{u}_{h}^{m}\right\rVert_{L^{2}}^{2}\right)^{\frac{1}{2}}\left\lVert\mathcal{P}_{\mathbb{U}_{h}}v\right\rVert_{L^{2}(0,T;L^{\infty}(\Omega))}$
@@
where we used Morrey’s inequality, the Sobolev embedding theorem and the stability of the $L^{2}$-projection operator, equation (3.7a), for the last inequalities. We proceed to term III:
$|\text{III}|$ $\leq\mu\left|\sum_{m=1}^{N}\int_{t^{m}}^{t^{m+1}}(\nabla\widetilde{u}^{m+1}_{h},\nabla\mathcal{P}_{\mathbb{U}_{h}}v(t-\Delta t))dt\right|$
$\leq\mu\left(\Delta t\sum_{m=1}^{N-1}\left\lVert\nabla\widetilde{u}^{m+1}_{h}\right\rVert_{L^{2}}^{2}\right)^{\frac{1}{2}}\left\lVert\nabla\mathcal{P}_{\mathbb{U}_{h}}v\right\rVert_{L^{2}([0,T]\times\Omega)}$
@@
$\leq\frac{C(E_{h}^{0}+\left\lVert f\right\rVert_{L^{2}([0,T]\times\Omega)}^{2})^{1/2}h^{s-1}}{\sqrt{\Delta t}}\left\lVert v\right\rVert_{L^{2}(0,T;H^{s}(\Omega))},$
@@
$|\text{V}|\leq\left|\int_{\Delta t}^{T+\Delta t}(f_{h},v(t-\Delta t))dt\right|\leq\left\lVert f_{h}\right\rVert_{L^{2}(0,T;L^{2}(\Omega))}\left\lVert v\right\rVert_{L^{2}(0,T;L^{2}(\Omega))}\leq C\left\lVert f\right\rVert_{L^{2}(0,T;L^{2}(\Omega))}\left\lVert v\right\rVert_{L^{2}(0,T;L^{2}(\Omega))}.$
@@
$\sup_{0\neq v\in L^{2}(0,T;H^{1}_{\operatorname{div}}\cap H^{s})}\left|\frac{(\partial_{t}\widehat{u}_{h},v)}{\left\lVert v\right\rVert_{L^{2}(0,T;H^{1}_{\operatorname{div}}\cap H^{s})}}\right|\leq C\left(1+\frac{h^{s-1}}{\sqrt{\Delta t}}\right),$
which is uniformly bounded under the assumption that $h^{s-1}\leq C\sqrt{\Delta t}$. Thus $\partial_{t}\widehat{u}_{h}\subset L^{2}(0,T;(H^{1}_{\operatorname{div}}(\Omega)\cap H^{s}(\Omega))^{*})$ uniformly in $h>0$.
Next, we will show that the sequences $\{\widetilde{u}_{h}\}_{h>0}$, $\{\widehat{u}_{h}\}_{h>0}$, $\{u_{h}\}_{h>0}$, $\{\bar{u}_{h}\}_{h>0}$ have the same weak limit $\widetilde{u}=\widehat{u}=u=\bar{u}$:
@@
Assume that $\Delta t=o_{h\to 0}(1)$. Then, we have $u=\bar{u}=\widetilde{u}=\widehat{u}$ a.e. in $0,T]\times\Omega$ and in $L^{2}([0,T]\times\Omega)$, and
(4.11) $\lim_{h\to 0}\left\lVert u_{h}-\widehat{u}_{h}\right\rVert_{L^{2}([0,T]\times\Omega)}$ $=0,$
(4.12) $\lim_{h\to 0}\left\lVert\widehat{u}_{h}-\widetilde{u}_{h}\right\rVert_{L^{2}([0,T]\times\Omega)}$ $=0,$
@@
$\left\lVert\bar{u}_{h}-\widetilde{u}_{h}\right\rVert_{L^{2}([0,T]\times\Omega)}^{2}$ $=\Delta t\sum_{m=0}^{N-1}\int_{\Omega}\left|\widetilde{u}^{m}_{h}-\widetilde{u}^{m+1}_{h}\right|^{2}dx$
$\leq 2\Delta t\sum_{m=0}^{N-1}\int_{\Omega}\left|u^{m}_{h}-\widetilde{u}^{m+1}_{h}\right|^{2}+\left|\widetilde{u}_{h}^{m}-u_{h}^{m}\right|^{2}dx$
@@
by the energy balance, Lemma 3.4 and Remark 3.5. This shows (4.13). Using this and that both $\bar{u}_{h}$ and $\widetilde{u}_{h}$ converge weakly, it then follows easily that for any test function $\varphi\in L^{2}([0,T]\times\Omega)$,
@@
$\left\lVert\widehat{u}_{h}-u_{h}\right\rVert_{L^{2}([0,T]\times\Omega)}^{2}$ $=\sum_{m=0}^{N-1}\int_{t^{m}}^{t^{m+1}}\!\!\int_{\Omega}\left|\frac{t-t^{m}}{\Delta t}(u^{m+1}_{h}-\widetilde{u}^{m+2}_{h})+\frac{t^{m+1}-t}{\Delta t}(u^{m}_{h}-\widetilde{u}^{m+1}_{h})\right|^{2}dxdt$
$\leq 2\Delta t\sum_{m=0}^{N-1}\int_{\Omega}\left|u^{m+1}_{h}-\widetilde{u}^{m+2}_{h}\right|^{2}+\left|\widetilde{u}^{m+1}_{h}-u^{m}_{h}\right|^{2}dx$
@@
Thus as $\Delta t,h\to 0$, we obtain (4.11), and in the same way as before, one shows that the limits $u$ and $\widehat{u}$ agree almost everywhere. Lastly, we have
$\left\lVert\widehat{u}_{h}-\widetilde{u}_{h}\right\rVert_{L^{2}([0,T]\times\Omega)}^{2}$ $=\sum_{m=0}^{N-1}\int_{t^{m}}^{t^{m+1}}\!\!\int_{\Omega}\left|\frac{t-t^{m}}{\Delta t}(\widetilde{u}^{m+2}_{h}-\widetilde{u}^{m+1}_{h})\right|^{2}dxdt$
$\leq\Delta t\sum_{m=0}^{N-1}\int_{\Omega}\left|\widetilde{u}^{m+2}_{h}-\widetilde{u}^{m+1}_{h}\right|^{2}dx$
$\leq 2\Delta t\sum_{m=0}^{N-1}\int_{\Omega}\left|u^{m+1}_{h}-\widetilde{u}^{m+2}_{h}\right|^{2}+\left|\widetilde{u}^{m+1}_{h}-u^{m+1}_{h}\right|^{2}dx$
which implies (4.12) and in the same way as before, that the limits $\widetilde{u}$ and $\widehat{u}$ agree almost everywhere. ∎
As a next step, we show that the limit $u$ is weakly divergence free, i.e., $(u(t),\nabla q)=0$ for almost every $t\in[0,T]$ and $q\in H^{1}(\Omega)$. Since $u\in L^{2}(0,T;H^{1}_{0}(\Omega))$ this then implies after integration by parts (the boundary term vanishes) that $\operatorname{div}u=0$ a.e.. This follows from the next lemma:
@@
Any weak limit $u$ of the sequence $\{u_{h}\}_{h>0}$ satisfies for a.e. $t\in[0,T]$,
$(u,\nabla q)=0,\quad\forall\,q\in H^{1}(\Omega).$
Proof.
Since $u$ is the weak limit of $\{{u}_{h}\}_{h>0}$, we have (up to a subsequence)
$(u,\nabla q)=\lim_{h\to 0}({u}_{h},\nabla q),$
for any $q\in H^{1}(\Omega)$ for a.e. $t\in[0,T]$. Then since ${u}_{h}(t)\in H^{1}_{0}(\Omega)$ for any fixed $h>0$ and $t\in[0,T]$, we can integrate by parts in the scheme (3.12b) to obtain
(4.14) $0=(\operatorname{div}{u}_{h},q)=-({u}_{h},\nabla q),$
for $q\in\mathbb{P}_{h}$. Then we can write
$|({u}_{h},\nabla q)|=|({u}_{h},\nabla(q-\mathcal{P}_{\mathbb{P}_{h}}q))|\leq\left\lVert{u}_{h}\right\rVert_{L^{2}}\left\lVert\nabla(q-\mathcal{P}_{\mathbb{P}_{h}}q)\right\rVert_{L^{2}}\stackrel{{\scriptstyle h\to 0}}{{\longrightarrow}}0,$
since ${u}_{h}\in L^{\infty}([0,T];L^{2}(\Omega))$ uniformly in $h>0$ and $\mathcal{P}_{\mathbb{P}_{h}}q\to q$ in $H^{1}(\Omega)$ as $h\to 0$ for any $q\in H^{1}(\Omega)$ by (3.7). Thus, in the limit $(u,\nabla q)=0$ for any $q\in H^{1}(\Omega)$ and a.e. $t\in[0,T]$. ∎
Next, we would like to apply the Aubin-Lions-Simon lemma [26] and obtain a strongly convergent subsequence of $\{u_{h}\}_{h>0}$ in $L^{2}([0,T]\times\Omega)$. However, we have the time continuity for $\widetilde{u}_{h}$ with values in the dual space of $H^{1}_{\operatorname{div}}(\Omega)\cap H^{s}(\Omega)$ and the spatial regularity in $H^{1}_{0}(\Omega)$ for the sequence of approximations $\{\widetilde{u}_{h}\}_{h>0}$, but they are not divergence free. In order to apply the Aubin-Lions lemma, we would either need them to be divergence free or uniform bounds $\partial_{t}\widetilde{u}_{h}\in L^{2}(0,T;(H^{1}_{0}(\Omega)\cap H^{s}(\Omega))^{*})$ for some $s\in\mathbb{N}$. However, this does not seem possible to achieve due to the lack of pressure bounds. We therefore need a variant of a classical lemma by Lions [29, Ch.III, Lem. 2.1], also called Ehrling’s lemma, adapted to our setting. The lemma we need is inspired by a similar result in [12]. We first define the space
$(\mathcal{P}_{\mathbb{V}_{h}}v,w)=(v,w),\quad\forall w\in\mathbb{V}_{h}.$
thus
We note that due to (4.14), we have $\mathcal{P}_{\mathbb{V}_{h}}\widetilde{u}^{m+1}_{h}=u^{m+1}_{h}$ for all $m$ and hence $u_{h}=\mathcal{P}_{\mathbb{V}_{h}}\widehat{u}_{h}(\cdot-\Delta t)$.
Lemma 4.4.
We have that
Proof.
implying the claim. ∎